A 5.53E+1 g sample of water at 9.967E+1 °C is placed in a constant pressure calorimeter. Then, 2.33E+1 g of zinc metal at 2.17E+1 °C is added to the water and the temperature drops to 9.63E+1 °C. What is the specific heat capacity of the zinc metal measured in this experiment?
Answer #1

Q = mc∆T

Q = heat energy (Joules, J), m = mass of a substance (g)

c = specific heat (units J/g∙oC), is a symbol meaning "the change in"

∆T = change in temperature (oC Celcius)

The heat lost from water = The heat gained by Zinc

For Water

m = 55.3 gm   c = 4.184 J/g∙oC ∆T = 99.67 oC - 96.3 oC = 3.37 oC

For Zinc

m = 23.3 gm   c = ? J/g∙oC ∆T = 96.3 oC - 21.7 oC = 74.6 oC

55.3 x 4.184 J/g∙oC x 3.37 oC = 23.3 gm x Sp ht of Zn x  74.6 oC

Sp ht of Zn = 779.73 Joules / 1738.18 g∙oC = 0.44859 J/g∙oC

specific heat capacity of the zinc metal measured in this experiment is 0.44859 J/g∙oC